OpenStudy (tiffani):

How do I solve this equation? 25(x+2)^2-2=0

OpenStudy (anonymous):

get the (x+2)^2 by itself first! :)

OpenStudy (anonymous):

1)Add 2 on both sides 2)Divide by 25 on both sides

OpenStudy (tiffani):

Doing it right now

OpenStudy (tiffani):

(x+2)^2=2/25?

OpenStudy (saifoo.khan):

Take a square root on both sides, \[(x+2) = \sqrt{ \frac{2}{25}}\]

OpenStudy (saifoo.khan):

\[x+2 = \frac{\sqrt2}{\sqrt{25}}\]

OpenStudy (saifoo.khan):

\[x+2 = \pm \frac{\sqrt2}{5}\]

OpenStudy (tiffani):

\[x+2=\sqrt{2}/5\]

OpenStudy (saifoo.khan):

dont forget the positive and negative sign.

OpenStudy (tiffani):

oh ya, now how do I get x by itself?

OpenStudy (saifoo.khan):

use a calc, \(x = \frac{\sqrt2}{5} - 2\) and \(x = -\frac{\sqrt2}{5} - 2\)

OpenStudy (tiffani):

Thanks

OpenStudy (saifoo.khan):

Welcome.

OpenStudy (tiffani):

would \[-2+-\sqrt{2}/5 \] be the same?

OpenStudy (saifoo.khan):

Nope, we will get 2 different answers.

OpenStudy (tiffani):

would \[2+-\sqrt{2}/5\]be?

OpenStudy (saifoo.khan):

\[-2+(-\frac{\sqrt2}{5}) = 1.7\]

OpenStudy (saifoo.khan):

And \[+2+\frac{\sqrt2}{5} = 2.28\]

OpenStudy (anonymous):

\[2 \pm \frac{\sqrt{2}}{5} ?\]

OpenStudy (anonymous):

that form?

OpenStudy (tiffani):

yes:)

OpenStudy (anonymous):

by the way that is -2 not positive 2

OpenStudy (anonymous):

\[-2 \pm \frac{\sqrt{2}}{5}\]

OpenStudy (anonymous):

\[x+2 = \pm \frac{\sqrt{2}}{5}\] Subtract 2 on both sides \[x=-2 \pm \frac{\sqrt{2}}{5}\]

OpenStudy (anonymous):

25[x2+4x+4]-2=0 25x2+100x+100-2=0 25x2+100x+98=0 using formula,solve for quadratic equation though Saifoo's method is better...